# On the complexity of matrix Putinar’s Positivstellensatz

### v1

@@ 5.3. Convergence rate of matrix SOS relaxations
where $f(x)\in\mathbb{R}[x]$, $G(x)$ is an $m\times m$ symmetric polynomial matrix. Denote by $f_{\min}$ the optimal value of (5.11).
A standard approach for solving (5.11) globally is the sum-of-squares hierarchy introduced in [15, 25, 43], consisting of a sequence of semidefinite programming relaxations. Recall that for a degree $k$, the $k$th degree truncation of $\mbox{QM}[G]$ is
$\mbox{QM}[G]_{k}:=\left\{\begin{array}[]{l|l}\sigma+\sum_{i=1}^{t}v_{i}^{T}Gv_{i}&\begin{array}[]{l}\sigma\in\Sigma[x],~{}v_{i}\in\mathbb{R}[x]^{m},\\ \deg(\sigma)\leq k,~{}\deg(v_{i}^{T}Gv_{i})\leq k\end{array}\end{array}\right\}.$
@@
Let $f_{k}$ denote the optimal value of (5.12). It holds the monotonicity relation
$f_{k}\leq f_{k+1}\leq\cdots\leq f_{\min}.$
The hierarchy (5.12) is said to have asymptotic convergence if $f_{k}\rightarrow\infty$ as $k\rightarrow\infty$ and it has finite convergence if $f_{k}=f_{\min}$ for all $k$ big enough. Under the archimedeanness, it was shown in [15, 25, 43] that the hierarchy (5.12) has asymptotic convergence. Recently, Huang and Nie [22] proved that this hierarchy also has finite convergence if in addition the nondegeneracy condition, strict complementarity condition and second order sufficient condition hold at every global minimizer of (5.11). In the following, we provide a polynomial bound on the convergence rate of the hierarchy (5.12).
@@
where $\kappa$, $\eta$ are respectively Lojasiewicz constant and exponent as in (5.3) and $C>0$ is a constant independent of $f$, $G$.
Proof.
@@
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{3\|f\|_{B}}{\varepsilon}\right)^{7\eta+3}.\$
@@
6. Conclusions and discussions
Under the archimedeanness, we prove a polynomial bound on degrees of terms appearing in matrix Putinar’s Positivstellensatz. For the special case that the constraining polynomial matrix is diagonal, it improves the exponential bound shown in [14]. As a byproduct, we also obtain a polynomial bound on the convergence rate of matrix SOS relaxations, which settles an open question posed by Dinh and Pham [10]. When the constraining set is unbounded, a similar bound is provided for matrix Putinar–Vasilescu’s Positivstellensatz.
Theorem 1.1 uses the Lojasiewicz inequality (5.3) for equivalent scalar polynomial inequalities of $K$. There also exists the Lojasiewicz inequality for polynomial matrices (see [9, 10]), whose Lojasiewicz exponent is typically smaller than the scalar one as in (5.3). An interesting future work is to give a proof based on the Lojasiewicz inequality for $G(x)$ directly and it may provide a better bound. On the another hand, the polynomial matrix $F(x)\succeq 0$ on the set $K$ described by $G(x)\succeq 0$ is equivalent to that $F(x)\succeq 0$ on the quantifier set
$\{x\in\mathbb{R}^{n}\mid y^{T}G(x)y\geq 0,~{}\forall~{}\|y\|_{2}=1\}.$
It would also be interesting to study the complexity of Putinar type Positivstellensatze with universal quantifiers for the above reformulation. We refer to [18] for the recent work on Positivstellensatze with universal quantifiers.
When $K$ is given by scalar polynomial inequalities, it was shown in [2, 3] that the Lojasiewicz exponent $\eta$ as in (3.5) is equal to $1$ if the standard linear independence constraint qualification holds at every $x\in K$. Then, we can get a representation with order $\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{10}$. An analogous of the linear independence constraint qualification in nonlinear semidefinite programming is the nondegeneracy condition [47, 51]. It is also interesting to investigate if there exist similar results for polynomial matrix inequalities under the nondegeneracy condition.
Acknowledgements: The author would like to thank Prof. Konrad Schmüdgen for helpful discussions on Proposition 4.1, Prof. Tien-Son Pham for fruitful suggestions and comments which help improve this paper.

### v2

@@ 5.3 Convergence rate of matrix SOS relaxations
where $f(x)\in\mathbb{R}[x]$, $G(x)$ is an $m\times m$ symmetric polynomial matrix. We denote the optimal value of (5.11) as $f_{\min}$.
A standard approach for solving (5.11) globally is the sum-of-squares hierarchy introduced in [16, 28, 48]. This hierarchy consists of a sequence of semidefinite programming relaxations. Recall that for a degree $k$, the $k$th degree truncation of $\mbox{QM}[G]$ is
$\mbox{QM}[G]_{k}:=\left\{\begin{array}[]{l|l}\sigma+\sum_{i=1}^{t}v_{i}^{T}Gv_{i}&\begin{array}[]{l}\sigma\in\Sigma[x],~{}v_{i}\in\mathbb{R}[x]^{m},~{}t\in\mathbb{N},\\ \deg(\sigma)\leq k,~{}\deg(v_{i}^{T}Gv_{i})\leq k\end{array}\end{array}\right\}.$
@@
Let $f_{k}$ denote the optimal value of (5.12). It holds the monotonicity relation: $f_{k}\leq f_{k+1}\leq\cdots\leq f_{\min}.$ The hierarchy (5.12) is said to have asymptotic convergence if $f_{k}\rightarrow\infty$ as $k\rightarrow\infty$, and it is said to have finite convergence if $f_{k}=f_{\min}$ for all $k$ big enough. When the quadratic module $\mbox{QM}[G]$ is Archimedean, it was shown in [16, 28, 48] that the hierarchy (5.12) converges asymptotically. Recently, Huang and Nie [24] proved that this hierarchy also has finite convergence if, in addition, the nondegeneracy condition, strict complementarity condition and second order sufficient condition hold at every global minimizer of (5.11). In the following, we provide a polynomial bound on the convergence rate of the hierarchy (5.12).
@@
where $\kappa$, $\eta$ are respectively the Łojasiewicz constant and exponent in (5.3), and $C>0$ is a constant independent of $f$, $G$. Furthermore, the Łojasiewicz exponent can be estimated as
$\eta\leq(3^{m-1}d_{G}+1)(3^{m}d_{G})^{n+\theta(m)-2}.$
Proof 5.8.
@@
$k\geq C\cdot{8}^{7\eta}\cdot 3^{6(m-1)}\theta(m)^{3}n^{2}d_{G}^{6}\kappa^{7}d^{14\eta}\left(\frac{3\|f\|_{B}}{\varepsilon}\right)^{7\eta+3},\$
@@
6 Conclusions and discussions
In this paper, we establish a polynomial bound on the degrees of terms appearing in matrix Putinar’s Positivstellensätz. For the special case that the constraining polynomial matrix is diagonal, this result improves the exponential bound shown in [15]. As a byproduct, we also obtain a polynomial bound on the convergence rate of matrix SOS relaxations, which resolves an open question posed by Dinh and Pham [10]. When the constraining set is unbounded, a similar bound is provided for matrix Putinar–Vasilescu’s Positivstellensätz.
Theorem 1.1 uses the Łojasiewicz inequality (5.3) for equivalent scalar polynomial inequalities of $K$. There also exists the Łojasiewicz inequality for polynomial matrices (see [9, 10]), where the corresponding Łojasiewicz exponent is typically smaller than the scalar one as in (5.3). An interesting future work is to provide a proof based directly on the Łojasiewicz inequality for $G(x)$, which may yields a better bound. On the another hand, the set $\{x\in\mathbb{R}^{n}\mid G(x)\succeq 0\}$ can be equivalently described by the quantifier set $\{x\in\mathbb{R}^{n}\mid y^{T}G(x)y\geq 0,~{}\forall~{}\|y\|_{2}=1\}.$ It would also be interesting to study the complexity of Putinar-type Positivstellensätze with universal quantifiers for the above reformulation. We refer to [19] for the recent work on Positivstellensätze with universal quantifiers.
When $K$ is given by scalar polynomial inequalities, it was shown in [45] that the Łojasiewicz exponent $\eta$ as in (3.5) is equal to $1$ if the linear independence constraint qualification holds at every $x\in K$. Consequently, we can get a representation with order $\left(\frac{\|F\|_{B}}{F_{\min}}\right)^{10}$. An analogue to the linear independence constraint qualification in nonlinear semidefinite programming is the nondegeneracy condition [52, 56]. It would also be interesting to explore whether similar results exist for polynomial matrix inequalities under the nondegeneracy condition.
Acknowledgements: The author would like to thank Prof. Konrad Schmüdgen for helpful discussions on Proposition 1, Prof. Tien-Son Pham for fruitful comments on this paper. The author also thanks the associate editor and three anonymous referees for valuable comments and suggestions.